OpenStudy (anonymous):

Translate into a mathematical statement 1) A number increased by 7 is 36.

OpenStudy (anonymous):

let x be the number, therefore x + 7 = 36

OpenStudy (anonymous):

Get it? Simple once you know.

OpenStudy (anonymous):

x = 36 - 7 x = 29